How do you foil complex numbers with conjugate that have been added to a variable?! Ex.
(x-7-5i)(x-7+5i)
Please do this step by step...the answer is x^2-14x+74 WHERE DOES THAT COME FROM? Ive been working on learning this for hours and I cant find it anywhere! Thanks!

Answer by Alan3354(69443)

How do you foil complex numbers with conjugate that have been added to a variable?! Ex.
(x-7-5i)(x-7+5i)
Please do this step by step...the answer is x^2-14x+74 WHERE DOES THAT COME FROM? Ive been working on learning this for hours and I cant find it anywhere!
--------------
Why do expect to find it somewhere? You have to multiply it.
----
(x-7-5i)(x-7+5i) = ((x+7) - 5i)*((x+7) + 5i)
= (x+7)^2 - (5i)^2
= x^2 + 14x + 49 + 25
= x^2 + 14x + 74